OpenStudy (sleepyjess):

What logarithmic function represents the data in the table? x f(x) 216 3 1296 4 7776 5

OpenStudy (sleepyjess):

Typing the equations now

OpenStudy (sleepyjess):

\[f(x)=\log_{x}5\] \[f(x)=6 logx\] \[f(x)=\log6^x\] \[f(x)=\log _{6}x\]

OpenStudy (amoodarya):

216=6^3 1296=6*216=6^4 7776 =6*1296=6^5

OpenStudy (amoodarya):

so f(X)=log (base=6) x

OpenStudy (sleepyjess):

Ok, thank you. I think I was just getting the 6 and x confused on where they should go because I ad originally thought it was C.